OpenStudy (anonymous):

An angle whose supplement is 3 times the measure of it supplement

OpenStudy (anonymous):

you dont have a complete sentence

OpenStudy (anonymous):

\[180-\theta=3(180-\theta) \implies \theta=180 degrees\]

OpenStudy (anonymous):

the english is screwy

OpenStudy (anonymous):

i did what I understood ;)

OpenStudy (anonymous):

well that is how the question is wrote out for me?

OpenStudy (anonymous):

you got an absurd conclusion

OpenStudy (anonymous):

no , write it again and double check

OpenStudy (anonymous):

Find the measure of an angle whose supplement is 3 times the angle.

OpenStudy (anonymous):

Thats the question

OpenStudy (anonymous):

oops, the last word is complement, sorry its late

OpenStudy (anonymous):

x + y = 180, y = 3x

OpenStudy (anonymous):

\[180-\theta=3(90-\theta) \implies 2\theta=90 \implies \theta=45\]

OpenStudy (anonymous):

3x + x = 180 x = 45

OpenStudy (anonymous):

y = 135

OpenStudy (anonymous):

\[if \ \theta>90\ then\ 180-\theta=3(\theta-90) \implies \theta=112.5\ else\ \theta=45\]

OpenStudy (anonymous):

I'm not sure what ask but i think 2 way can solve frist: 180-x=3(90-x) or: 180=3(90-x) 180-x=270-3x 180=270-3x -x+3x=270-180 -270+180=-3x 2x = 90 -90 = -3x x= 90 -90/-3=x x=45 30 = x

OpenStudy (anonymous):

@Nancy Lam: you are assuming that x<90. What if x>90?